OpenStudy (anonymous):

Find a formula for the function f that is defined by y= f(x) <-> 2x+ 2xy+ y^2=5 and y> -x

OpenStudy (loser66):

What does <-> mean?

OpenStudy (anonymous):

What are you looking for when you say "formula?" What mathematics course are you taking. That will help narrow it down.

OpenStudy (swissgirl):

<-> = if and only if @Loser66

OpenStudy (anonymous):

I think it just means y= f(x) equals that equation. I just have to write a formula using the notation f(x) =something. Calculus and Analytical Geometry, we just work with functions that have one variable

OpenStudy (anonymous):

So i think I need to get it in terms of y

OpenStudy (anonymous):

Should I use the quadratic formula?

OpenStudy (swissgirl):

Thats what i was thinking

OpenStudy (anonymous):

Think of the left side as a quadratic.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So I have to get it in the form ax^2+bx+c=0. Would I divide the entire equation by 5?

OpenStudy (swissgirl):

yes move over the 5 so you would have y^2+2xy+(2x-5)=0 where 2x-5=C

OpenStudy (swissgirl):

Nahhh you would subtract 5 from both sides 5-5=0

OpenStudy (anonymous):

Think of "2x" as a constant. Factor the left side accordingly.

OpenStudy (anonymous):

So would 2x be like the coefficient? when I'm trying to figure out what a, b, and c are to solve for quadratic formula

OpenStudy (swissgirl):

a=1 b=2x and c=2x-5

OpenStudy (anonymous):

Oh. Thank you. So how am I supposed to give medals to you guys. What do I click?

OpenStudy (anonymous):

Appreciate the help. All of you

OpenStudy (anonymous):

No problem. Did the quadratic formula work out yet?

OpenStudy (anonymous):

I think I got it from here. The rest is easier algebra

OpenStudy (anonymous):

\[x= -x \pm(\sqrt{4x^2-8x+20}/2)\]

OpenStudy (anonymous):

Okie doke. Not sure about the medals. I'm new here. Happy mathing!

OpenStudy (anonymous):

\[-2x +\sqrt{-6x-20} / 2\]

OpenStudy (anonymous):

Make sure you divide both the 2x and the sqrt by two.

OpenStudy (anonymous):

\[\sqrt{b^2-4(a)(c)} = \sqrt{(2x)^2-4(1)(2x-5)} = \sqrt{4x^2 - 8x - 20}\]

OpenStudy (anonymous):

My final answer: \[x \pm \sqrt{x^2-2x+5}\] as your two factors. You would have to multiply them to get your equation for f(x).

OpenStudy (anonymous):

So.... \[(x +\sqrt{x^2-2x+5})(x-\sqrt{x^2-2x+5})\]